# A complete proof of the geometric Bombieri–Lang conjecture for ramified covers of abelian varieties

### v1

@@ 1.2 Ideas of the proof
There are three difficulties in realizing the intuitive idea above. Firstly, $({\mathcal{D}}_{\rm red})_{t_{\infty}}$ might be non-reduced. In this case the tangency of $({\mathcal{D}}_{\rm red})_{t_{\infty}}$ and the entire curve in ${\mathcal{A}}\times\{t_{\infty}\}$ is trivial. In order to overcome this problem, we prove the following (slightly surprising) lemma in §3.2.
@@
The second difficulty is that $\mathcal{D}$} may contain the whole fibre ${\mathcal{A}}\times\{t_{\infty}\}$. In this case the tangency of $({\mathcal{D}}_{\rm red})_{t_{\infty}}$ and the entire curve in ${\mathcal{A}}\times\{t_{\infty}\}$ is also trivial. To overcome this problem, we use a weak semistable reduction theorem (cf. [AK, Theorem 0.3]). Namely, there is a generically finite proper surjective map $T^{\prime}\to T$ and a proper birational equidimensional morphism ${\mathcal{X}}^{\prime}\to{\mathcal{X}}\times_{T}T^{\prime}$ such that the fibre of ${\mathcal{X}}^{\prime}\to T^{\prime}$ are all reduced. This reducedness essentially implies that the ramification divisor of $X^{\prime}\to A\times_{K}T^{\prime}_{K}$ is “horizontal” over $T^{\prime}_{K}$, where $X^{\prime}$ is the generic fibre of the composition ${\mathcal{X}}^{\prime}\to{\mathcal{X}}\times_{T}T^{\prime}\to{\mathcal{X}}\to B$. This resolves the second difficulty.
@@
To overcome this problem, we will use the canonical model $X_{\rm can}$ of $X$ in the minimal model program, which is a birational model of $X$ with ample $\mathbb{Q}$-Cartier canonical class. Moreover, it has mild singularities. By the Hacon–Mckernan extension theorem (cf. [HM, Corollary 1.7]), we have a birational morphism $X_{\rm can}\to X$ whenever $X$ does not contain rational curves. We can use this canonical model to avoid the case that $R$ is not $\mathbb{Q}$-Cartier or $K_{X}$ has base locus. See §3.4 for more details.
The above is the rough idea of the proof, but of course it is definitely not rigorous. For example, the meaning of “an excessive number of” points is not clear. We will use Nevanlinna theory (see §3.1) to make the proof rigorous. Specifically, we will employ the counting function in Nevanlinna theory to quantify the number of intersections between $x_{n}(B)$ and $\mathcal{R}$}, as well as the number of tangent points between $(s_{n},t_{n})(B)$ and ${\mathcal{D}}_{\rm red}$. Eventually, we will show two things: on the one hand, the positivity of $R$ implies that the tangent points between $(s_{n},t_{n})(B)$ and ${\mathcal{D}}_{\rm red}$ account for a positive proportion of all intersection points (in the asymptotic sense; see (3.4.14) and Remark 3.18); on the other hand, by Yamanoi’s Second Main Theorem (see Theorem 3.6(1)) and some limit arguments (Lemma 1.7 mentioned above and Lemma 3.5), we will prove that this proportion of tangent points cannot be positive (see Lemma 3.19), thereby derive a contradiction.
Unlike the classical Nevanlinna theory, we need to consider the counting functions (as well as proximity and characteristic functions) for holomorphic maps from an open disk to a complex variety, rather than only from the whole complex plane. Furthermore, we need to vary the holomorphic maps and study the limit behaviour of these three functions in Nevanlinna theory (see Lemma 3.5).
@@ 1.3 Notations and terminology
We denote $\mathbb{N}$} to be the set of non-negative integers and $\mathbb{N}$$^{+}$ to be the set of positive integers.
@@
We use $\sim$ to denote the linear equivalent relation between $\mathbb{Q}$-Weil, $\mathbb{Q}$-Cartier divisors. We use $\cong$ to denote the isomorphism between two $\mathbb{Q}$-line bundles.
For a proper morphism $f\colon X\to Y$ between $K$-varieties, there is a push-forward homomorphism $f_{*}\colon Z_{k}X\to Z_{k}Y$ between $k$-cycles. More precisely, for any closed subvariety $V$ of $X$, the image $W=f(V)$ is a closed subvariety of $Y$. If $\dim W<\dim V$, set $f_{*}[V]=0$. If $\dim W=\dim V$, then the function field $K(V)$ of $V$ is a finite extension of the function field $K(W)$ of $W$. Set $f_{*}[V]=[K(V):K(W)]\cdot[W]$. This extends linearly to a homomorphism $f_{*}\colon Z_{k}X\to Z_{k}Y$ for each $k\geq 0$. By [Ful, Theorem 1.4], $f_{*}$ descends to a homomorphism $f_{*}\colon A_{k}(X)\to A_{k}(Y)$ between the groups of $k$-cycles modulo rational equivalence. In particular, if $\dim X=\dim Y=d$ and k=d1k=d-1, $f$ induces a push-forward homomorphism $f_{*}\colon{\rm Cl}(X)\to{\rm Cl}(Y)$ between (Weil) divisor class groups.
@@
Let $K$ be a finitely generated field over a field $k$, and assume that $k$ is algebraically closed in $K$. Let $A$ be an abelian variety over $K$. Denote by $A^{(K/k)}$ Chow’s $K/k$-trace of $A$ over $k$. Denote
$V(A,K)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}(A(K)/A^{(K/k)}(k))\otimes_{\mathbb{Z}}\mathbb{R},$
which is a finite-dimensional $\mathbb{R}$-vector space. In our setting, $k$ is usually $\mathbb{C}$.
For a complex abelian variety $A$, the complex Lie algebra ${\rm Lie}(A)$ is defined to be the group of translation-invariant holomorphic derivations on ${\mathcal{O}}_{A}$. By restriction, we have canonical isomorphisms ${\rm Lie}(A)\cong T_{x}A$ for any $x\in A$.
For a variety $X$ over $\mathbb{C}$, denote by $TX\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}{\rm Spec}({\rm Sym}^{*}(\Omega_{X}))$ the (geometric) tangent bundle of $X$. The complex point of $TX$ is represented by a pair $(x,v)$ with $x\in X$ and $v\in T_{x}X$ a tangent vector of $X$ at $x$. For a morphism $f\colon X\to Y$ between complex varieties, there is a natural tangent map ${\rm d}f\colon TX\to TY$ induced by $f$, which is algebraic. For $x\in X$, the restriction of ${\rm d}f$ to $T_{x}X\subseteq TX$ coincides with the tangent map mentioned above.
For a normal variety $X$ over a field $k$, define the canonical sheaf $\omega_{X}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}i_{*}\omega_{U}$, where $U$ is the smooth locus of $X$ and $i\colon U\hookrightarrow X$ is the natural open immersion. This is a reflexive sheaf of rank 1, so it corresponds to a Weil divisor class $K_{X}$ of $X$. We call $K_{X}$ the canonical divisor of $X$.
@@
Let $X$ be a complex projective variety. Denote $X(\mathbb{C})$ its associated complex analytic space (in particular, $X(\mathbb{C})$ is endowed with Euclidean topology). Let $S$ be a subset of $X(\mathbb{C})$. Denote by $\overline{S}$ the closure of $S$ in $X(\mathbb{C})$ and by $\overline{S}^{\rm zar}$ the Zariski closure of $S$ in $X$. Assume that the smooth locus $X^{\rm sm}(\mathbb{C})$ is covered by countably many open balls $\{D_{\alpha}\}_{\alpha\in I}$. We say that $S$ is measurable if $S\cap D_{\alpha}$ is measurable under the Lebesgue measure of the ball $D_{\alpha}$ for all $\alpha$. We say that $S$ has measure 0 if $S\cap D_{\alpha}$ has measure 0 for all $\alpha$. Throughout this article, when we say “for almost all points in $X(\mathbb{C})$”, we always mean that for points of $X(\mathbb{C})$ outside a subset of measure 0.
For any positive real number $r$, denote by
${\mathbb{D}}_{r}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{z\in\mathbb{C}\big||z|<r\},\qquad\overline{{\mathbb{D}}}_{r}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{z\in\mathbb{C}\big||z|\leq r\}.$
the discs of radius $r$. Write ${\mathbb{D}}={\mathbb{D}}_{1}$ and $\overline{{\mathbb{D}}}=\overline{{\mathbb{D}}}_{1}$. Denote by $v_{\rm st}=\frac{\rm d}{{\rm d}z}$ the tangent vector of ${\mathbb{D}}_{r}$ at 0 under the standard coordinate $z$.
Let $(Y,d)$ be a metric space. Let $\{r_{n}\}_{n\geq 1}$ be a sequence of positive real numbers converging to infinity. Let $\phi_{n}\colon{\mathbb{D}}_{r_{n}}\to Y$ be a sequence of continuous maps. We say that $\{\phi_{n}\}_{n\geq 1}$ uniformly converges to a map $\phi\colon\mathbb{C}\to Y$ if for every $\epsilon>0$, there is $N\geq 1$ such that for every $n\geq N$ and $z\in{\mathbb{D}}_{r_{n}}$, $d(\phi(z),\phi_{n}(z))\leq\epsilon$. If such $\phi$ exists, it is unique and continuous. Moreover, if $Y$ is further a complex analytic variety and $\phi_{n}$ are holomorphic, then $\phi$ is holomorphic.
@@
Acknowledgements. I would like to express my sincere thanks to my advisors Junyi Xie and Xinyi Yuan for their valuable insights shared during discussions and for their consistent encouragement throughout the research. I am grateful to She Yang for many useful discussions in Lemma 3.7, and to Jihao Liu for providing the reference [AK]. I am appreciate to Ariyan Javanpeykar for many useful communications on the first version of this article. I would also like to thank Ya Deng, Ziyang Gao, Carlo Gasbarri, Katsutoshi Yamanoi and Umberto Zannier for numerous helpful comments on this work.
@@ 2 Locations of limit points
In this section we recall some results from [XY2], which will be used in the proof of the main result. §2.1 is some basic on the linear entire curves (cf. [XY2, §2]) and §2.2 explicitly describes the location of limit points of a given sequence of sections of an abelian scheme, which is related to the linear entire curves (cf. [XY2, §3]).
@@ 2.1 Linear entire curves
Linear entire curves
Let $A$ be a complex abelian variety. Let $x\in A$ be a point, and $v\in{\rm Lie}(A)$ be a nonzero vector in the Lie algebra. By definition, $v$ defines a translation-invariant vector field on $\mathcal{A}$}, and thus it is a foliation ${\mathcal{F}}(v)$ on $A$. The leaf ${\mathcal{F}}(v)_{x}={\mathcal{F}}(x,v)$ of this foliation through $x$ is called the linear entire curve through $x$ in the direction $v$. By translation, we have
@@
The linear entire curve is generally not algebraic in $A$. Its Zariski closure is equal to $A(v)+x$ for an abelian subvariety $A(v)$ of $A$. In fact, by translation, it suffices to assume $x=0$. Then $\phi_{(0,v)}$ is a homomorphism, so its image ${\mathcal{F}}(0,v)$ is actually a compact Lie group, and thus its Zariski closure $A(v)$ inherits a group structure. Moreover, $A(v)$ is irreducible as an analytic variety, since the image of $\mathbb{C}$ (under $\phi_{(0,v)}$) is irreducible. Then $A(v)$ is an abelian subvariety of $A$.
@@
The main result of this subsection is to give a clear description of Zariski closures of linear leaves in fibres of abelian schemes.
Let $B$ be a smooth quasi-projective cure over $\mathbb{C}$, and let $\pi\colon{\mathcal{A}}\to B$ be an abelian scheme. Let $K=\mathbb{C}(B)$ be the function field of $B$, and $A$ be the generic fibre of $\pi\colon{\mathcal{A}}\to B$. By the Lang–Néron theorem (cf. [XY1, Theorem 3.4]),
@@
Moreover, we have the following key lemma.
Lemma 2.1 ([XY2, Lemma 2.4]).
Let $H_{1},H_{2}$ be two abelian subvarieties of $A$, and let $H$ be the identity component of $H_{1}\cap H_{2}$. Then $V(H_{1},K)\cap V(H_{2},K)=V(H,K)$.
As a consequence, for every $s\in V(A,K)$, there exists a unique minimal abelian subvariety $G(s)$ of $A$ with $s\in V(G(s),K)$.
@@
Here the transfer map (associated to $s$) $\delta(s,\cdot)\colon T_{b}B\to{\rm Lie}({\mathcal{A}}_{b})$ is defined in [XY2, §2.2] using the Betti map. A key property of this transfer map is the following non-degeneracy theorem.
Theorem 2.3 ([XY2, Theorem 2.2]).
$(B,\pi\colon{\mathcal{A}}\to B,K,A)$ are as above. Let $s\in V(A,K)$ be a nonzero element. Let $\Sigma(s)$ be the set of $b\in B$ such that the transfer map
@@ 2.2 Locations of limit points
Theorem 2.4 ([XY2, Theorem 3.1(a)]).
@@
$(B,\pi\colon{\mathcal{A}}\to B,K,A)$ are as above. Let $\{s_{n}\}_{n\geq 1}$ be a sequence in ${\mathcal{A}}(B)$ satisfying the condition $(2)$ in Theorem 2.4. Denote by $G$ the smallest abelian subvariety of $A$ such that $s_{\infty}$ lies in $G(K)_{\mathbb{R}}$. Denote by $\mathcal{G}$} the Zariski closure of $G$ in $\mathcal{A}$}. Assume that the natural map ${\rm End}_{K}(A)\to{\rm{End}}_{\overline{K}}(A_{\overline{K}})$ is an isomorphism. Then for almost all $b\in B$, there is a linear entire curve in ${\mathcal{A}}_{b}$ which is the unique limit of a re-parametrization of a subsequence ((depending on $b)$ of $\{s_{n}\}_{n\geq 1}$, and the Zariski closure of this linear entire curve in ${\mathcal{A}}_{b}$ is a translation of ${\mathcal{G}}_{b}$.
@@
By Proposition 2.2 and Theorem 2.3, there are two subsets $S(A),\Sigma(s_{\infty})$ of $B$ of measure 0 such that for all $b\in B\setminus(S(A)\cup\Sigma(s_{\infty}))$, the Zariski closure of the linear leaf ${\mathcal{F}}(0,\delta(s_{\infty},v_{b}))$ in ${\mathcal{A}}_{b}$ is equal to ${\mathcal{G}}_{b}$, and the transfer map $\delta(s_{\infty},\cdot)\colon T_{b}(B)\to{\rm Lie}({\mathcal{A}}_{b})$ is nonzero. Fix one such $b$. By compactness, there is a limit point $x\in{\mathcal{A}}_{b}$ of the sequence $\{s_{n}(b)\}_{n\geq 1}$. Passing to a subsequence $\{n_{j}\}_{j\geq 1}\subseteq{\mathbb{N}},n_{j}\to\infty$, taking $b_{n_{j}}=b$ and $x=\lim_{j\to\infty}s_{n_{j}}(b)$ in Theorem 2.4, we get a linear entire curve $\phi_{(x,\delta(s_{\infty},v_{b}))}$ of ${\mathcal{A}}_{b}$ which is the unique limit of a re-parametrization of $\{s_{n_{j}}\}_{j\geq 1}$, and its Zariski closure in ${\mathcal{A}}_{b}$ is equal to the translation $x+{\mathcal{G}}_{b}$. This finishes the proof. ∎
Using the same method as in [XY2, Theorem 3.3], we can prove the following corollary which will be used in §3.4.

### v2

@@ 1.4 Three technical difficulties
Firstly, $({\mathcal{D}}_{\rm red})_{t_{\infty}}$ might be non-reduced. In this case the tangency of $({\mathcal{D}}_{\rm red})_{t_{\infty}}$ and the entire curve in ${\mathcal{A}}\times\{t_{\infty}\}$ is trivial. In order to overcome this problem, we prove the following (slightly surprising) lemma in §3.2.
@@
Non-flatness of $\mathcal{D}$} over $T$
The second difficulty is that $\mathcal{D}$} may contain the whole fibre ${\mathcal{A}}\times\{t_{\infty}\}$. In this case the tangency of $({\mathcal{D}}_{\rm red})_{t_{\infty}}$ and the entire curve in ${\mathcal{A}}\times\{t_{\infty}\}$ is also trivial. To overcome this problem, we use a weak semistable reduction theorem (cf. [AK00, Theorem 0.3]). Namely, there is a generically finite proper surjective map $T^{\prime}\to T$ and a proper birational equidimensional morphism ${\mathcal{X}}^{\prime}\to{\mathcal{X}}\times_{T}T^{\prime}$ such that the fibre of ${\mathcal{X}}^{\prime}\to T^{\prime}$ are all reduced. This reducedness essentially implies that the ramification divisor of $X^{\prime}\to A\times_{K}T^{\prime}_{K}$ is “horizontal” over $T^{\prime}_{K}$, where $X^{\prime}$ is the generic fibre of the composition ${\mathcal{X}}^{\prime}\to{\mathcal{X}}\times_{T}T^{\prime}\to{\mathcal{X}}\to B$. This resolves the second difficulty.
We emphasize that the essential reason why this operation transforms the problem is that $t_{n}\in T(\mathbb{C})$ is constant for each $n\geq 1$. Therefore the section $(s_{n},t_{n})\colon B\to{\mathcal{A}}\times_{\mathbb{C}}T$ can be lifted to a section $(s_{n},t^{\prime}_{n})\colon B\to{\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$. This is not always possible when $t_{n}\in T_{K}(K)$ is non-constant.
Problem caused by singularities
@@
To overcome this problem, we will use the canonical model $X_{\rm can}$ of $X$ in the minimal model program, which is a birational model of $X$ with ample $\mathbb{Q}$-Cartier canonical class. Moreover, it has mild singularities. By the Hacon–Mckernan extension theorem (cf. [HM07, Corollary 1.7]), we have a birational morphism $X_{\rm can}\to X$ whenever $X$ does not contain rational curves. We can use this canonical model to avoid the case that $R$ is not $\mathbb{Q}$-Cartier or $K_{X}$ has base locus. See §3.3 for more details.
@@ 1.5 Notations and terminology
We use $\sim$ to denote the linear equivalent relation between $\mathbb{Q}$-Weil or $\mathbb{Q}$-Cartier divisors.
@@
For a normal variety $X$ over a field $k$, define the canonical sheaf $\omega_{X}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}i_{*}\omega_{U}$, where $U$ is the smooth locus of $X$ and $i\colon U\hookrightarrow X$ is the natural open immersion. This is a reflexive sheaf of rank 1, so it corresponds to a Weil divisor class $K_{X}$ of $X$. We call $K_{X}$ a canonical divisor of $X$.
@@
For a variety $X$ over $\mathbb{C}$, denote $X(\mathbb{C})$ its associated complex analytic space (endowed with Euclidean topology). For a point $x\in X(\mathbb{C})$, denote $T_{x}X$ for the tangent space $T_{x}(X(\mathbb{C}))$ for simplicity. Denote by $TX\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}{\rm Spec}({\rm Sym}^{*}(\Omega_{X}))$ the (geometric) tangent bundle of $X$. The complex point of $TX$ is represented by a pair $(x,v)$ with $x\in X(\mathbb{C})$ and $v\in T_{x}X$ a tangent vector of $X$ at $x$. For a morphism $f\colon X\to Y$ between complex varieties, there is a natural tangent map ${\rm d}f\colon TX\to TY$ induced by $f$, which is algebraic. For $x\in X$, the restriction of ${\rm d}f$ to $T_{x}X\subseteq TX$ coincides with the tangent map mentioned above.
For any real number $r>0$, denote by ${\mathbb{D}}_{r}\subseteq\mathbb{C}$ the open disc of radius $r$ centred at 0. Let $(Y,d)$ be a metric space. Let $\{r_{n}\}_{n\geq 1}$ be a sequence of positive real numbers tending to infinity. Let $\phi_{n}\colon{\mathbb{D}}_{r_{n}}\to Y$ be a sequence of continuous maps. We say that $\{\phi_{n}\}_{n\geq 1}$ converges to a map $\phi\colon\mathbb{C}\to Y$ uniformly on any compact subset of $\mathbb{C}$ if for every $\varepsilon>0$ and every compact subset $\Omega\subseteq\mathbb{C}$, there is $N\geq 1$ such that for every $n\geq N$ and $z\in\Omega$, $d(\phi(z),\phi_{n}(z))\leq\varepsilon$. If such $\phi$ exists, it is unique and continuous. Moreover, if $Y$ is further a complex analytic variety and $\phi_{n}$ are holomorphic, then $\phi$ is holomorphic.
Acknowledgements. I would like to express my sincere thanks to my advisors Junyi Xie and Xinyi Yuan for their valuable insights shared during discussions and for their consistent encouragement throughout the research. I am grateful to She Yang for many useful discussions in Lemma 3.7, and to Jihao Liu for providing the reference [AK00]. I am appreciate to Ariyan Javanpeykar for many useful communications on the first version of this article. I would also like to thank Ya Deng, Ziyang Gao, Carlo Gasbarri, Katsutoshi Yamanoi and Umberto Zannier for numerous helpful comments on this work.
@@ 2 Locations of limit points
In this section we recall some results from [35], which will be used in the next section. §2.1 is some basic on the linear entire curves (cf. [35, §2]) and §2.2 explicitly describes the location of limit points of a given sequence of sections of an abelian scheme, which is related to the linear entire curves (cf. [35, §3]).
@@ 2.1 Linear entire curves
Let $A$ be a complex abelian variety. Denote by ${\rm Lie}(A)$ the group of translation-invariant holomorphic derivations on ${\mathcal{O}}_{A}$. Let $x\in A$ be a point, and $v\in{\rm Lie}(A)$ be a nonzero vector in the Lie algebra. By restriction, we have a canonical isomorphism ${\rm Lie}(A)\simeq T_{x}A$. By definition, $v$ defines a translation-invariant vector field on $\mathcal{A}$}, and thus it is a foliation ${\mathcal{F}}(v)$ on $A$. The leaf ${\mathcal{F}}(v)_{x}={\mathcal{F}}(x,v)$ of this foliation through $x$ is called the linear entire curve through $x$ in the direction $v$. By translation, we have
@@
Linear entire curves are generally not algebraic. We recall a clear description in [35] of Zariski closures of linear leaves in fibres of abelian schemes.
Let $B$ be a smooth quasi-projective cure over $\mathbb{C}$, and let $\pi\colon{\mathcal{A}}\to B$ be an abelian scheme. Let $K=\mathbb{C}(B)$ be the function field of $B$, and $A$ be the generic fibre of $\pi\colon{\mathcal{A}}\to B$. By the Lang–Néron theorem (cf. [CON06, Theorem 2.1]),
@@
Moreover, Xie–Yuan proved in [35, Lemma 2.4] that for every $s\in V(A,K)$, there exists a unique minimal abelian subvariety $G(s)$ of $A$ with $s\in V(G(s),K)$.
@@
Here the transfer map (associated to $s$) $\delta(s,\cdot)\colon T_{b}B\to{\rm Lie}({\mathcal{A}}_{b})$ is defined in [35, §2.2] using the Betti map. A key property of this transfer map is the following non-degeneracy theorem.
Theorem 2.2 ([35, Theorem 2.2]).
$(B,\pi\colon{\mathcal{A}}\to B,K,A)$ are as above. Let $s\in V(A,K)$ be a nonzero element. Let $\Sigma(s)$ be the set of $b\in B(\mathbb{C})$ such that the transfer map
@@ 2.2 Locations of limit points
For a real number $r>0$, denote by ${\mathbb{D}}_{r}\subseteq\mathbb{C}$ the open disc of radius $r$ centred at 0. Write ${\mathbb{D}}={\mathbb{D}}_{1}$. Denote by $v_{\rm st}=\frac{\rm d}{{\rm d}z}$ the tangent vector of ${\mathbb{D}}_{r}$ at 0 under the standard coordinate $z$.
Theorem 2.3 ([35, Theorem 3.1(a)]).
@@
$(B,\pi\colon{\mathcal{A}}\to B,K,A)$ are as above. Let $\{s_{n}\}_{n\geq 1}$ be a sequence in ${\mathcal{A}}(B)$ satisfying the condition $(2)$ in Theorem 2.3. Denote by $G$ the smallest abelian subvariety of $A$ such that $s_{\infty}$ lies in $G(K)_{\mathbb{R}}$. Denote by $\mathcal{G}$} the Zariski closure of $G$ in $\mathcal{A}$}. Assume that the natural map ${\rm End}_{K}(A)\to{\rm{End}}_{\overline{K}}(A_{\overline{K}})$ is an isomorphism. Then for all $b\in B(\mathbb{C})$ outside a subset of measure 0, there is a linear entire curve in ${\mathcal{A}}_{b}$ which is the unique limit of a re-parametrization of a subsequence ((depending on $b)$ of $\{s_{n}\}_{n\geq 1}$, and the Zariski closure of this linear entire curve in ${\mathcal{A}}_{b}$ is a translation of ${\mathcal{G}}_{b}$.
@@
By Proposition 2.1 and Theorem 2.2, there are two subsets $S(A),\Sigma(s_{\infty})$ of $B$ of measure 0 such that for all $b\in B(\mathbb{C})\setminus(S(A)\cup\Sigma(s_{\infty}))$, the Zariski closure of the linear leaf ${\mathcal{F}}(0,\delta(s_{\infty},v_{b}))$ in ${\mathcal{A}}_{b}$ is equal to ${\mathcal{G}}_{b}$, and the transfer map $\delta(s_{\infty},\cdot)\colon T_{b}(B)\to{\rm Lie}({\mathcal{A}}_{b})$ is nonzero. Fix one such $b$. By compactness, there is a limit point $x\in{\mathcal{A}}_{b}$ of the sequence $\{s_{n}(b)\}_{n\geq 1}$. Passing to a subsequence $\{n_{j}\}_{j\geq 1}\subseteq{\mathbb{Z}}_{>0}$, taking $b_{n_{j}}=b$ and $x=\lim_{j\to\infty}s_{n_{j}}(b)$ in Theorem 2.3, we get a linear entire curve $\phi_{(x,\delta(s_{\infty},v_{b}))}$ of ${\mathcal{A}}_{b}$ which is the unique limit of a re-parametrization of $\{s_{n_{j}}\}_{j\geq 1}$, and its Zariski closure in ${\mathcal{A}}_{b}$ is equal to the translation $x+{\mathcal{G}}_{b}$. This finishes the proof. ∎
Using the same method as in [35, Theorem 3.3], we can prove the following corollary which will be used in §3.4.